twice the difference between a number and twenty-five is three times the number. Find the number. How do you set up the problem and what is the answer. Thank you

Answer by ewatrrr(24785)


Hi,
Letting n represent the number, then we preceed to "set up the problem"
by writing as we read.
2(n-25)= 3n
Simplifying and Solving
2n - 50 = 3n
-50= n
checking our answer
2n - 50 = 3n
-100-50 = -150= 3*-50